OpenStudy (anonymous):

compute the difference quotient of the given function Q(t)=4-7t

OpenStudy (anonymous):

difference quotient is : \[\huge \frac{Q(t+h)-Q(t)}{h} \] so, give me the expression for Q(t+h)....

OpenStudy (anonymous):

f(x+h)=4-7(x+h)=4-7x+7h

OpenStudy (anonymous):

ok, it should be 4 - 7x - 7h, right? also, let's stick with the variable t: \[\large \huge \frac{Q(t+h)-Q(t)}{h}=\frac{(4-7t-7h)-(4-7t)}{h} \] can you simplify the top?

OpenStudy (anonymous):

yea... 4-7x+7h-(4-7t)

OpenStudy (anonymous):

is that the top? shouldn't it be -7h?

OpenStudy (anonymous):

yesss and yeah thats what I meant

OpenStudy (anonymous):

\(\large \frac{Q(t+h)-Q(t)}{h}=\frac{(4-7t-7h)-(4-7t)}{h}=\frac{4-7t-7h-4+7t}{h} \) \(\large =\frac{\cancel4\cancel{-7t}-7h\cancel{-4} \cancel{+7t}}{h}=\frac{-7h}{h} \) \)

OpenStudy (anonymous):

=....???

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

is it -7??

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

and, btw, i usually follow up on the question i help with but if by chance i forget, you can always tag the helper by putting a "@" sign and their name... this will call them back to the question... like this: @jessiedeee

OpenStudy (anonymous):

ok thankyou @dpaInc